# KAD: No More FAD! An Effective and Efficient Evaluation Metric for Audio Generation

### v1

@@ 2.1 Fréchet Audio Distance and its Limitations
    Normality assumption: The assumption that audio embeddings follow a Gaussian distribution often fails for real-world data. Such data are often asymmetrically distributed or unevenly clustered, which compromises the metric’s ability to accurately measure similarity. Similar limitations have been observed in the computer vision domain, where the normality assumption was shown to be unsuitable for the Inception embedding space used for FID [2]. Figure 2 shows how the actual distribution of VGGish embeddings from the Clotho dataset (reduced via UMAP) differs significantly from the samples drawn from a Gaussian distribution with the same mean and covariance. This deviation highlights the limitations of the normality assumption for representing real-world audio data, leading to potential inaccuracies and over- or under-estimations in FAD values.
@@ 2.2 Maximum Mean Discrepancy
Several previous studies have explored MMD-based metrics for evaluating generative models such as the Kernel Inception Distance (KID) proposed by Binkowski et al. [39]. KID is computed as the MMD between samples embedded by the Inception model [40], yielding an unbiased measure with rapidly decaying deviation as sample size increases – addressing some of the well-known limitations of FID. However, Inception-based scores including FID have been shown to diverge considerably from human evaluations [41]. In response, Jayasumana et al. [2] introduced the CLIP Maximum Mean Discrepancy (CMMD), which leverages the CLIP [42] image encoder to generate embeddings that are more human-aligned when assessing generative models.
@@ 3 Kernel Audio Distance
where $\alpha$ is a resolution scaling factor introduced for convenient score comparison. We set $\alpha=1000$ as the default.
@@ 3.2 Kernel Function and Bandwidth Selection
KAD relies on evaluating pairwise relationships between embeddings through a kernel function. Many commonly used kernels such as Gaussian, Laplacian, and Matérn kernels are examples of a characteristic kernel, meaning that the MMD it induces i) fully distinguishes between two embedding distributions and ii) is zero if and only if the two distributions under test are identical [43]. This property is crucial for model evaluation, as it ensures that the KAD metric captures meaningful differences in the embedding distributions of real and generated audio samples. As an example, a second-order polynomial kernel $(1+x^{T}y)^{2}$ cannot differentiate between distributions with the same mean and variance, but different kurtosis [44].
@@ 4.1 Experiment 1: Reliability of KAD in Perceptual Alignment
We compute both KAD and FAD using embedding from several well-known models, including VGGish [7], PANNs [64], CLAP [65], PaSST [66], and OpenL3 [67], all of which are trained on environmental sounds. These embedding models are widely used for the calculation of FAD scores for text-to-audio [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and vision-to-audio generation[21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Since music-focused models can differ substantially in their learned representations, we also include MERT [68] and CLAP-laion-music [69] for completeness. We then measure the Spearman rank correlation between each metric’s scores and the average human evaluation scores, and also the p-value. Correlations with $p>0.05$ are shaded in 3 to indicate a lack of statistical significance.
@@
Among the tested embedding models, PANNs-WGLM(WaveGram-LogMel) achieves the strongest correlation with human judegments, aligning with prior research that highlighted its suitability for FAD-based evaluations [4]. Based on this observation, we select PANNs-WGLM as the primary embedding model in subsequent experiments to further investigate the performance of KAD.
@@ 4.3 Experiment 3: Computation Cost Comparison
For the measurements, AMD EPYC 7413 CPU (24 cores) and an Nvidia RTX 3090 GPU were used, and the code is implemented on PyTorch for both KAD and FAD calculations. For FAD, we refactored the Microsoft FAD toolkit [5] for consistency in CPU/GPU usage, thereby ensuring comparability of runtime measurements. All values were calculated in single-precision floating points.
@@ Appendix A Median Pair-wise Distance Heuristic for the Kernel Bandwidth
We compared the MMD scores between the clean audio files from the Clotho dataset’s eval split, and the same files under various audio signal transformations, including Gaussian noise injection, bit depth reduction, low-pass filtering, dynamic range reduction (limiter), and pitch shifts. The degradations were applied using the audiomentations python library [71].
@@ Appendix B Bias of FAD
We release a toolkit named kadtk that could calculate KAD scores from input audios. Given input audio directories for the ground-truth reference set and target evaluation set, the toolkit calculates the score and saves it to the given output filepath in csv format. The toolkit is written in Python and supports Pytorch and Tensorflow environments (refer to our Readme document for further details). It supports numerous models for embedding extraction: CLAP [65, 69], Encodec [72], MERT [68], VGGish [7], PANNs [64], OpenL3 [67], PaSST [66], DAC [73], CDPAM [74], Wav2vec2.0 [75], HuBERT [76], WavLM [77], and Whisper [78]. This covers a wide range of the audio domain including general environment sounds (sound effects, foley sounds, etc.), music, and speech. We also support FAD calculation for comparison.
The kadtk is released under the MIT License, allowing unrestricted use, modification, and distribution with proper attribution. The full license text is included in the repository. Some codes were brought from the FAD toolkit (fadtk) [5, 79, 80]. We sincerely thank the authors for sharing the code as open source. Note that fadtk was also licensed under the MIT License.

### v2

@@ 2.1 Fréchet Audio Distance and its Limitations
    Normality assumption: The assumption that audio embeddings follow a Gaussian distribution often fails for real-world data. Such data are often asymmetrically distributed or unevenly clustered, which compromises the metric’s ability to accurately measure similarity. Similar limitations have been observed in the computer vision domain, where the normality assumption was shown to be unsuitable for the Inception [37] embedding space used for FID [2]. Figure 2 shows how the actual distribution of VGGish embeddings from the Clotho dataset (reduced via UMAP) differs significantly from the samples drawn from a Gaussian distribution with the same mean and covariance. This deviation highlights the limitations of the normality assumption for representing real-world audio data, leading to potential inaccuracies and over- or under-estimations in FAD values.
@@ 2.2 Maximum Mean Discrepancy
One of the most widely used MMD-based metrics for evaluating generative models is the Kernel Inception Distance (KID) proposed by Bińkowski et al. [40], as the squared MMD value between Inception embeddings of images with a cubic polynomial kernel. KID was used as an evaluation metric in several audio-related works [41, 42, 43, 44], particularly in music generation. For image quality, Jayasumana et al. [2] proposed the CLIP Maximum Mean Discrepancy (CMMD) for evaluation on CLIP embeddings [45] and a Gaussian kernel, with Novack et al. [46] following similar practices but on CLAP [47] embeddings. Many of these works acknowledge the potential advantages of MMD. However, to our knowledge, there has been no comprehensive study of its reliability in comparison to FAD or the effect of various design choices.
@@ 3 Kernel Audio Distance
where $\alpha$ is a resolution scaling factor introduced for convenient score comparison. We set $\alpha=100$ as the default.
@@ 3.2 Kernel Function and Bandwidth Selection
KAD relies on evaluating pairwise relationships between embeddings through a kernel function. Many commonly used kernels such as Gaussian, Laplacian, and Matérn kernels are examples of a characteristic kernel, meaning that the MMD it induces i) fully distinguishes between two embedding distributions and ii) is zero if and only if the two distributions under test are identical [48]. This property is crucial for model evaluation as it ensures that the KAD metric captures meaningful differences in the embedding distributions of real and generated audio samples. As an example, a cubic polynomial kernel $(x^{T}y+1)^{2}$ cannot differentiate between distributions with the same mean, variance and skewness, but different kurtosis [49].
@@ 4.1 Experiment 1: Reliability of KAD in Perceptual Alignment
We compute both KAD and FAD using embedding from several well-known models, including VGGish [7], PANNs [69], CLAP [70, 47], and PaSST [71], all of which are trained on environmental sounds. These embedding models are widely used for the calculation of FAD scores for text-to-audio [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20] and vision-to-audio generation[21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Since music-focused models can differ substantially in their learned representations, we also include MERT [72] and CLAP-laion-music [47] for completeness. We then measure the Spearman rank correlation between each metric’s scores and the average human evaluation scores, as well as the p-value. Correlations with $p>0.05$ are shaded in 3 to indicate a lack of statistical significance.
@@
Among the tested embedding models, PANNs-WGLM(WaveGram-LogMel) achieves the strongest correlation with human judgments, aligning with prior research that highlighted its suitability for FAD-based evaluations [4]. Based on this observation, we select PANNs-WGLM as the primary embedding model in subsequent experiments to further investigate the performance of KAD.
@@ 4.3 Experiment 3: Computation Cost Comparison
For the measurements, AMD EPYC 7413 CPU (24 cores) and an Nvidia RTX 3090 GPU were used, and the code is implemented on PyTorch for both KAD and FAD calculations. For FAD, we refactored the Microsoft FAD toolkit [5] for consistency in CPU/GPU usage, thereby ensuring the comparability of runtime measurements. All values were calculated in single-precision floating points.
@@ Appendix A Median Pair-wise Distance Heuristic for the Kernel Bandwidth
We compared the MMD scores between the clean audio files from the Clotho dataset’s eval split, and the same files under various audio signal transformations including Gaussian noise injection, bit depth reduction, low-pass filtering, dynamic range reduction (limiter), and pitch shifts. The degradations were applied using the audiomentations python library [74].
@@ Appendix B Bias of FAD
We release a toolkit named kadtk that can calculate KAD scores from input audios. Given input audio directories for the ground-truth reference set and target evaluation set, the toolkit calculates the score and saves it to the given output file path in CSV format. The toolkit is written in Python and supports Pytorch and Tensorflow environments (refer to our Readme document for further details). It supports numerous models for embedding extraction: CLAP [70, 47], Encodec [75], MERT [72], VGGish [7], PANNs [69], OpenL3 [76], PaSST [71], DAC [77], CDPAM [78], Wav2vec2.0 [79], HuBERT [80], WavLM [81], and Whisper [82]. This covers a wide range of the audio domain including general environment sounds (sound effects, foley sounds, etc.), music, and speech. We also support FAD calculation for comparison.
The kadtk is released under the MIT License, allowing unrestricted use, modification, and distribution with proper attribution. The full license text is included in the repository. Some codes were brought from the FAD toolkit (fadtk) [5, 83, 84]. We sincerely thank the authors for sharing the code as an open source. Note that fadtk was also licensed under the MIT License.
For the computational cost comparison in Section 4.3, we used torch.linalg.eigval and torch.sqrt to compute the covariance matrix of FAD, ensuring a fair comparison with KAD in a fully parallelized GPU setting. However, this approach is prone to accuracy issues, often leading to discrepancies with fadtk, which relies on scipy.linalg.sqrtm (available only for CPU computation). To ensure high FAD score accuracy at the cost of increased runtime, we used scipy.linalg.sqrtm for the perceptual alignment experiment (Section 4.1) and in the released toolkit.